OpenStudy (anonymous):

Write the standard equation for the circle that passes through the points : (1,1) , (1,3) , (9,2) PLEASE HELP (most of the work done)

OpenStudy (anonymous):

I found that E= -4, that D= (-79/8) and that F=(95/8) The answer is suppost to be 8x^2 + 8y^2 ...... =0 Instead of x^2 + y^2 + Dx + Ey + F = 0

OpenStudy (lastdaywork):

Actually, your answer (x^2 + y^2 + Dx + Ey + F = 0) also correctly represents the circle. But if you specifically want it in the format (8x^2 + 8y^2 ...... =0) then simply multiply each term by 8.

OpenStudy (anonymous):

so..... 8x^2 + 8y^2 + (-79x) + (-32y) + 95 = 0 ?

OpenStudy (lastdaywork):

Yep!

OpenStudy (anonymous):

Thanks, you're a bro.

OpenStudy (lastdaywork):

You're welcome :)